# Query of CC: Unearthing Large Scale Domain-Specific Knowledge from Public Corpora

### 26 Jan 2024 v1

@@ Abstract
Zhaoye Fei1,2*, Yunfan Shao1,2♠, Linyang Li1,2♠, Zhiyuan Zeng1,2 Hang Yan2, Xipeng Qiu1 and Dahua Lin2 ${}^{1}$chool of Computer Science, Fudan University, Shanghai, China ${}^{2}$hanghai AI Laboratory {zyfei20, yfshao19, xpqiu}@fudan.edu.com {lilinyang, yanhang}@pjlab.org.cn
**footnotetext: Work done during internship at Shanghai AI Laboratory.footnotetext: These authors contributed equally to this work.
@@ 4 Knowledge Pile
Leveraging Query of CC, Based on queries of some knowledge categories, we retrieved high-quality knowledge data from processed public corpora. We call the collected datasets as Knowledge Pile. In this section, we will introduce the analysis of queries (Section 4.1) and the statistic information of Knowledge Pile (Section 4.2), See more implement details in Appendix A.
@@ 4.2 Data Statistics
www.semanticscholar.org 134252
www.ipl.org 83076
slideplayer.com 80583
www.ncbi.nlm.nih.gov 60623
www.bartleby.com 54546
link.springer.com 50807
pubmed.ncbi.nlm.nih.gov 50409
quizlet.com 43359
www.studymode.com 38685
www.innovations-report.com 37845
@@
Based on the proposed method, we have formed a high-quality knowledge dataset Knowledge Pile, which maintains about 900GB disk and 300B tokens (using Llama2 tokenizer). As shown in Figure 1, comparing with other datasets in academic and mathematical reasoning domains, we have acquired a large-scale, high-quality knowledge dataset at a lower cost, without the need for manual intervention. Through automated query bootstrapping, we efficiently capture the information about the seed query. Knowledge Pile not only covers mathematical reasoning data but also encompasses rich knowledge-oriented corpora spanning various fields such as biology, physics, etc., enhancing its comprehensive research and application potential.
Figure 4: Timestamp Statistics of Knowledge Pile
5 Experiment
Based on Knowledge Pile, we further-pretraining two models: Llama2-QoC and Mistral-QoC, which is built upon the foundations of Llama2 and Mistral. This section will present the details of our experiment, including model training (Section 5.1), baselines (Section 5.2), and evaluation (Section 5.3).
@@ 5 Experiment
5.1 Model Training Details
In this study, we employed the InternLM [^1] https://github.com/InternLM/InternLM InternLM-Team (2023) library for training all models on 256 A800 GPUs with bfloat16 mixed precision. Data parallelism was also utilized during the training process. We employed the AMSP Chen et al. (2023), shard optimizer state across 8 cards to reduce communication overhead. Simultaneously, data parallelism was employed for training. To enhance throughput and reduce memory consumption, we introduced the Flash attention 2 Dao (2023) module. All models underwent training for 50,000 steps, with a global batch size of 4 million tokens per step, totaling 200 billion training tokens.
5.2 Baselines
We compare the performance with some open source pre-trained models: Llama2 Touvron et al. (2023b), Code-Llama Rozière et al. (2023), Baichuan 2-Base Yang et al. (2023), Mistral Jiang et al. (2023), Qwen 2 Bai et al. (2023) and some language models for mathematical reasoning: Llemma Azerbayev et al. (2023) and Minerva Lewkowycz et al. (2022). It is noteworthy that Llemma is a model for mathematical reasoning based on the continued pre-training of Code Llama in the mathematical dataset Proof of Pile 2. Minerva is a large-scale mathematical model derived from the continued pre-training of PaLM.
5.3 Evaluation
During the evaluation, we utilized the open-source library OpenCompass  [^2] https://github.com/open-compass/opencompass, which serves as a platform for evaluating large models. OpenCompass offers various evaluation datasets and supports efficient task partitioning to maximize the utilization of computational resources. For the selection of evaluation datasets, we opted for three distinct capabilities to assess both Llama2-QoC and Mistral-QoC. These encompassed mathematical reasoning datasets such as Math, GSM8K, knowledge-oriented language understanding datasets including MMLU, AGIEval, and challenging reasoning tasks BIG-Bench hard. More details for evaluation will be described in App B.
6 Results and Discussion
6.1 Overview
@@ 7 Conclusion
In this study, we propose an efficient method Query of CC, for the automated collection of specialized domain data. Leveraging seed data from some specific domains, we employ a large language model for query bootstrapping. By optimizing the breadth and depth of queries, we expand the query to retrieve data relevant to the specified domain. Ultimately, we collected and released an open dataset comprising approximately 900GB of Knowledge Pile, equivalent to approximately 300 billion tokens, in the fields of mathematics and knowledge. Experimental results demonstrate that the adoption of Knowledge Pile significantly enhances the model’s performance in some reasoning tasks, such as math word problems and professional examinations. Our objective is not only to establish a research foundation for community studies in mathematical and knowledge-related reasoning but also to provide an efficient and cost-effective method for collecting high-quality data, thereby facilitating the accumulation of more high-quality data.
Appendix A Collection implement details
In the implement section, we primarily discuss three components, the public corpora (A.1), retrieval engine (A.2), and the post-processing settings (A.3).
A.1 Public Corpora
A.2 Retrieval Engine
A.3 Post Processing
Appendix B Evaluation Dataset Details

### 05 Mar 2024 v3

@@ Abstract
Zhaoye Fei1,2*, Yunfan Shao1,2♠, Linyang Li1,2♠, Zhiyuan Zeng1,2 Conghui He2, Hang Yan2, Dahua Lin2 and Xipeng Qiu1 ${}^{1}$chool of Computer Science, Fudan University, Shanghai, China ${}^{2}$hanghai AI Laboratory {zyfei20, yfshao19, xpqiu}@fudan.edu.cn {lilinyang, yanhang}@pjlab.org.cn
**footnotetext:  Work done during internship at Shanghai AI Laboratory.footnotetext:  These authors contributed equally to this work.11footnotetext:  Knowledge Pile will be released in https://huggingface.co/datasets/ngc7293/Knowledge_Pile.
@@ 4 Knowledge Pile
Leveraging Query of CC, Based on queries of some knowledge categories, we retrieved high-quality knowledge data from processed public corpora. We call the collected datasets as Knowledge Pile. In this section, we will introduce the analysis of queries (Section 4.1) and the statistic information of Knowledge Pile (Section 4.2), See more implement details in Appendix C.
@@ 4.2 Data Statistics
en.wikipedia.org 398833
www.semanticscholar.org 141268
slideplayer.com 108177
www.ncbi.nlm.nih.gov 97009
link.springer.com 85357
www.ipl.org 84084
pubmed.ncbi.nlm.nih.gov 68934
www.reference.com 61658
www.bartleby.com 60097
quizlet.com 56752
@@
Based on the proposed method, we have formed a high-quality knowledge dataset Knowledge Pile, which maintains about 735GB disk and 188B tokens (using Llama2 tokenizer). As shown in Figure 1, comparing with other datasets in academic and mathematical reasoning domains, we have acquired a large-scale, high-quality knowledge dataset at a lower cost, without the need for manual intervention. Through automated query bootstrapping, we efficiently capture the information about the seed query. Knowledge Pile not only covers mathematical reasoning data but also encompasses rich knowledge-oriented corpora spanning various fields such as biology, physics, etc., enhancing its comprehensive research and application potential.
Figure 4: Timestamp Statistics of Knowledge Pile
5 Experiment and Discussion
5.1 Implementation Details
Based on Knowledge Pile, we further train two models: Llama2-QoC and Mistral-QoC, based on Llama2 and Mistral.
Training
In our experiment, we employed the InternLM [^1] https://github.com/InternLM/InternLM InternLM-Team (2023) library for training all models on 256 A800 GPUs with bfloat16 mixed precision, and only utilized data parallelism during the training process. To enhance throughput and reduce memory consumption, we introduced the Flash attention 2 Dao (2023) module. Above these, both Llama-QoC and Mistral-QoC achieve 4000 tokens per GPU per second(TGS). More training details will be described in Appendix A.
Evaluation
During the evaluation, we utilized the open-source library OpenCompass  [^2] https://github.com/open-compass/opencompass, which serves as a platform for evaluating large models. Leveraging Opencompass, we compare the performance with some open source pre-trained models: Llama2 Touvron et al. (2023b), Code-Llama Rozière et al. (2023), Baichuan 2-Base Yang et al. (2023), Mistral Jiang et al. (2023), Qwen 2 Bai et al. (2023) and some language models for mathematical reasoning: Llemma Azerbayev et al. (2023) and Minerva Lewkowycz et al. (2022). For the selection of evaluation datasets, we opted for three distinct capabilities to assess both Llama2-QoC and Mistral-QoC. These encompassed mathematical reasoning datasets such as Math, GSM8K, knowledge-oriented language understanding datasets including MMLU, AGIEval, and challenging reasoning tasks BIG-Bench hard. More details for evaluation will be described in App B.
5.2 Main Results
Figure 5: The results in MMLU benchmark
@@ 6 Conclusion
In this study, we propose an efficient method Query of CC, for the automated collection of specialized domain data. Leveraging seed data from some specific domains, we employ a large language model for query bootstrapping. By optimizing the breadth and depth of queries, we expand the query to retrieve data relevant to the specified domain. Ultimately, we collected and released an open dataset comprising approximately 735GB of Knowledge Pile, equivalent to approximately 188 billion tokens, in the fields of mathematics and knowledge. Experimental results demonstrate that the adoption of Knowledge Pile significantly enhances the model’s performance in some reasoning tasks, such as math word problems and professional examinations. Our objective is not only to establish a research foundation for community studies in mathematical and knowledge-related reasoning but also to provide an efficient and cost-effective method for collecting high-quality data, thereby facilitating the accumulation of more high-quality data.
7 Limitation
In this paper, we propose a collection method for domain-specific data, Query of CC, which extends domain-relevant queries through LLM bootstrapping for data retrieval. Additionally, we demonstrate through empirical evaluation on further training of LLAMA and Mistral that data collected using this method significantly improves the ability of LLM in some specific domains. However, this method still has the following potential limitations:
Data Quality
The quality of data collected by Query of CC depends largely on the quality of data in public corpora. In this work, we utilized the Common Crawl corpus as our public corpus, extracting and processing the CC dump up to April 2023. Despite leveraging some methods Wenzek et al. (2020) for processing high-quality web data, we found that the corpus still contains an amount of low-quality and erroneously extracted content. Thus, improving the data quality of public corpora is also a direction for future research.
Data Scale
The data source of Query of CC is limited by public corpora, and unable to collect information not present in these corpora. This limitation requires particular attention. Moreover, the proportion of domain data in public databases also influences the scale of the collected dataset. During the collection process, scoring based on similarity scores can be employed to filter highly relevant data.
Appendix A Model Training Details
In this study, we train all models with bf16 mixed precision and only data parallel. We employed the AMSP Chen et al. (2023), shard optimizer state across 8 cards to reduce communication overhead. Simultaneously, data parallelism was employed for training. To enhance throughput and reduce memory consumption, we introduced the Flash attention 2 Dao (2023) module. All models underwent training for 50,000 steps, with a global batch size of 4 million tokens per step, totaling 200 billion training tokens.
Appendix B Evaluation
During the evaluation, we utilized the open-source library OpenCompass  [^3] https://github.com/open-compass/opencompass, which serves as a platform for evaluating large models. OpenCompass offers various evaluation datasets and supports efficient task partitioning to maximize the utilization of computational resources. For the selection of evaluation datasets, we opted for three distinct capabilities to assess both Llama2-QoC and Mistral-QoC. These encompassed mathematical reasoning datasets such as Math, GSM8K, knowledge-oriented language understanding datasets including MMLU, AGIEval, and challenging reasoning tasks BIG-Bench hard. The details of evaluation datasets are as follows:
@@ Appendix B Evaluation
Appendix C Collection implement details
In the implement section, we primarily discuss three components, the public corpora (C.1), retrieval engine (C.2), and the post-processing settings (C.3).
C.1 Public Corpora
C.2 Retrieval Engine
C.3 Post Processing